help please



U how you arrived 1. Simplify the following complex fractions: x 1 (10 points)
Answer #1

3x+1 2-32 3x+1 2-30 3x+1 2-3x 1 + 3 SONe 1 + 3 4+3 ya-9 ya = 9+3 y (443) y2-9 4 (9735 19+3)(y-3) = Y-30 (2 sove: Xog 22-X-6 x2+68-77 3x+2) - (x-3) (x+3) & x2-3x+28-6 x2+7X-X-7 3(x+7) = (x-3)(x+3) 212-3)+2131-3) X(x+7)-1(x+7) 3(= (2+3) (0-1) = 3(X+3) (X4) (x+2) = 3x+9 x2+286-x-2 - 3x+9 & +3-a Ans